A life insurance representative believes that the mean age of people who buy their first life insurance plan is less than 35. To test his belief he takes a random sample of 15 customers who have just purchased their first life insurance. Their ages are 42, 43,28, 34, 30, 36, 25, 29, 32, 33, 27, 30, 22, 37, and 40. There is not enough evidence to say the data are nonnormal. Can we conclude at the 1% significance level that the insurance representative is correct?
Answer #1

Solution:

x x2
42 1764
43 1849
28 784
34 1156
30 900
36 1296
25 625
29 841
32 1024
33 1089
27 729
30 900
22 484
37 1369
40 1600
--- ---
\sumx=488 \sumx2=16410

The sample mean is \bar x

Mean \bar x   = (\sumx / n) )

=42+43+28+34+30+36+25+29+32+33+27+30+22+37+40 /15

=488 /15

=32.5333

Mean \bar x   =32.53

The sample standard is S

  S =\sqrt{}( \sum x2 ) - (( \sum x)2 / n ) n -1


=\sqrt{}(16410-(488)215 )14

=\sqrt{}(16410-15876.2667)14

=\sqrt{}533.7333 /14

=\sqrt{}38.1238

=6.174

This is the left tailed test .

The null and alternative hypothesis is ,

H0 :  \mu  = 35

Ha : \mu    \neq 35

Test statistic = t

= (\bar x - \mu ) / s / \sqrt n

= (35 - 32.53) / 6.17 / \sqrt 15

= −1.550

Test statistic = t =−1.550

P-value =0.0717

\alpha = 0.01

P-value ≥ \alpha

0.0717 ≥ 0.01

Do not reject the null hypothesis .

There is insufficient evidence to suggest that